Act 3 of The 2003 Leonid meteor shower was expected to last 24 hours. The margin of error for this prediction was 0.7 days. What was the range for the length of time that Act 3 could last?
Answer by stanbon(75887)

Act 3 of The 2003 Leonid meteor shower was expected to last 24 hours. The margin of error for this prediction was 0.7 days. What was the range for the length of time that Act 3 could last?
----
Lower Limit: 24-0.7(24) = 0.3(24) = 7.2 hrs
---
Upper Limit: 24+0.7(24) = 40.8 hrs
=======================================
Cheers,
Stan H.
===========